OpenStudy (anonymous):

1/(sqrt(3){3}-sqrt(2))= 1/(3^4-2^4)

OpenStudy (unklerhaukus):

\[\frac{1}{\sqrt[3]{3}-\sqrt{2}}=\frac {1}{3^4-2^4}\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Try rationalizing... @hollywood_chrissy

OpenStudy (phi):

that is not true.

OpenStudy (anonymous):

Ok, so rationalizing will not work @phi ?

OpenStudy (phi):

the question is missing something. as posted, the left side ≠ the right side

OpenStudy (unklerhaukus):

so, false ?

OpenStudy (anonymous):

Yes! There is something missing...

OpenStudy (anonymous):

\[\large{\sqrt[3]{3} -\sqrt{2} \ne 3^4 - 2^4 }\]

OpenStudy (anonymous):

@Hollywood_chrissy please check your quest. again.